OpenStudy (anonymous):

I need to find the indefinite integral of 2/(1 + 4x^2). I have looked through every example in my textbook and every note from class, as well as asked another student in the class, and cannot find how to do this.

OpenStudy (anonymous):

look at arctan all you need to do to use arctan formula is substitution: 2x=u try it & let me know

OpenStudy (anonymous):

\[=\int\limits_{ }^{}2dx/(1+4x ^{2})=\int\limits_{}^{}du/(1+u ^{2})=\arctan u=\arctan (2x)+ const\]